# SynPlay: Importing Real-world Diversity for a Synthetic Human Dataset

### v1

@@ Abstract
We introduce Synthetic Playground (SynPlay), a new synthetic human dataset that aims to bring out the diversity of human appearance in the real world. We focus on two factors to achieve a level of diversity that has not yet been seen in previous works: i) realistic human motions and poses and ii) multiple camera viewpoints towards human instances. We first use a game engine and its library-provided elementary motions to create games where virtual players can take less-constrained and natural movements while following the game rules (i.e., rule-guided motion design as opposed to detail-guided design). We then augment the elementary motions with real human motions captured with a motion capture device. To render various human appearances in the games from multiple viewpoints, we use seven virtual cameras encompassing the ground and aerial views, capturing abundant aerial-vs-ground and dynamic-vs-static attributes of the scene. Through extensive and carefully-designed experiments, we show that using SynPlay in model training leads to enhanced accuracy over existing synthetic datasets for human detection and segmentation. The benefit of SynPlay becomes even greater for tasks in the data-scarce regime, such as few-shot and cross-domain learning tasks. These results clearly demonstrate that SynPlay can be used as an essential dataset with rich attributes of complex human appearances and poses suitable for model pretraining. SynPlay dataset comprising over 73k images and 6.5M human instances, is available for download at https://synplaydataset.github.io/.
[Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] [Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image] Figure 1: SynPlay dataset is constructed while players play six traditional games in a virtual playground also introduced in the Netflix TV show “Squid game” [24]. We have diversified the human appearances in the scenes by focusing on two factors: i) leveraging real-world human motions and ii) adopting multiple viewpoints.
@@ 1 Introduction
Large-scale synthetic datasets, known for their scalability, provide a practical solution to the increasing demand for training large-capacity models. Recently developed rendering engines (e.g., Unity [49], Unreal [16], etc.) have significantly enhanced the realism of synthetic data, broadening its applicability across various computer vision tasks. Despite efforts to scale synthetic data to match the extensive curation of real-world data, the desired level of diversity has not yet been achieved. This insufficiency in diversity is largely due to the inadequate consideration and integration of key factors that are essential to real-world diversity in the process of creating synthetic data.
Recently, several attempts have been made to increase human appearance diversity by controlling innate characteristics (e.g., race, gender) [6], body shape (e.g, height) [36, 6], or clothing [6]. These datasets have demonstrated their effectiveness in tasks aimed at identifying human characteristics from close-up images, e.g., human body/pose estimation [36, 6] and shape reconstruction [33]. However, these datasets have seldom yielded a discernible positive impact on computer vision tasks aimed at identifying humans from a distance, e.g., human detection and segmentation.
When it comes to recognizing the overall human appearance from a distance, the motions and poses exhibited by the individuals play more vital roles than other characteristics. Despite prior attempts to synthesize various human poses, the quality of the rendering remained suboptimal, lacking in realism [39] and diversity [45]. AMASS [33] was the output of an early endeavor aimed at achieving both realism and diversity, where a motion scanner was utilized to collect real human motions. Before capturing these motions, detailed descriptions were provided to articulate specific movements — e.g., 5 seconds waving above the head with both arms [^1] This motion description was used to construct the Mocap Database HDM05 in AMASS (https://resources.mpi-inf.mpg.de/HDM05/05-01/index.html), while adhering to physical constraints that limit large movements in motion-capture environments. This detail-guided motion design often results in capturing a restricted range of motions tied to specific descriptions while missing out on all the motions that defy easy description.
We claim that providing relatively high-level, less-detailed guidance greatly helps in breaking out from the aforementioned limitations and provides more freedom towards the expansion of the diversity in human motions. In constructing our dataset, we follow the new rule-guided motion design approach providing game “rules” or winning strategies for the virtual players to follow, which serve as a set of significantly coarser guidelines when compared to detail-guided approaches. In this way, the motions that they manifest are not confined to predetermined/easily-describable motions. As for the “rules,” we opted to borrow them from the six traditional Korean games that were also played in the Netflix TV series “Squid Game” [24]. These games involve substantial amounts of physical movements, which naturally provide room for a diverse set of human poses and motions. The diversity is further influenced by in-game factors such as the uniquely defined rules of each game, the number of players, and the interactions between them.
Under our rule-guided motion design approach, each scenario run (i.e., one round of a game played with specific settings) in the virtual environment is initialized by carrying out a scenario design step which is followed by the incorporation of real-world motions. The scenario design involves the setup of all the knobs that control the appearance, players (winners and losers), game dynamics (e.g., how/when each game ends), and the human motion evolutions for each specific scenario. This is where the high-level rules of a given game are defined, and the coarse boundary of how human motions can evolve within the game is set. The incorporation of real-world motion is the phase where a rich variety of motions truly comes to life. Details on the entire pipeline will be elaborated in Section 3.
In addition, we also took into account that human appearance can vary greatly depending on the perspective from which it is viewed. Accordingly, we capture every scene from multiple viewpoints by implementing several image-capturing devices to take advantage of different perspective-related characteristics: three Unmanned Aerial Vehicles (UAVs), three Closed-Circuit Televisions (CCTVs), and one Unmanned Ground Vehicle (UGV). The three UAVs fly with random trajectories at different altitudes, the three CCTVs are located at the front, side, and back of the game playground, and the UGV moves randomly within the playground where the game is being played. These devices offer a variety of image-capturing properties, including aerial-vs-ground and dynamic-vs-static. Our strategy, designed to provide very diverse viewpoints in scene capture process, not only serves to ensure that the dataset includes more diverse human appearances but also broadens the potential tasks (e.g., re-identification, multi-view applications, aerial-to-ground scene matching, etc.) for which the dataset can be used.
By leveraging the aforementioned human appearance-diversifying strategies, we construct a large-scale synthetic human dataset called SynPlay that contains more than 73k images with 6.5M human instances, see sample images in Fig 1. To demonstrate SynPlay’s ability to represent a variety of human appearances to the extent seen in the real world, we conduct a series of experiments where we evaluate the impact of leveraging SynPlay alongside real (non-synthetic) datasets curated for a variety of human-related tasks, i.e.,, aerial-view/ground-view human detection and segmentation. For all the tasks, training with SynPlay outperforms its counterparts (i.e., training-from-scratch, using other synthetic data) across a variety of datasets. Experiments also demonstrate that the SynPlay dataset significantly improves model performance on data-scarce tasks, highlighting its value in scenarios that require substantial supplementary training data.
@@ 2 Related Works
Synthetic human data. The creation of various synthetic human datasets has been facilitated by the advancement of modern synthetic data rendering engines such as Blender [7], Unity [49], and Unreal [16], alongside human modeling tools like MakeHuman [34] and Character Creator [47]. These rendering engines enable a realistic representation of humans in 3D virtual environments, while the modeling tools give creators precise control over the design of virtual characters. The creators of these datasets leveraged these tools to meticulously control key design factors, ensuring suitability for specific tasks, e.g., SOMAset [4], PersonX [48], UnrealPerson [52], CARGO [51] for Re-Identification, SURREAL [50] for pose estimation, GTA5 [39] for semantic segmentation, and Archangel-Synthetic [45] for detection.
Recently, several attempts have been made to enhance the realism of virtual human models, with the aim of bringing them closer to the quality of their real-world counterparts. SMPL-X [37] and AMASS [33] used motion capturing devices to capture natural human motions. BEDLAM [6] tried to improve the diversity of various factors such as skin-tones or clothing that affect human outer appearance, still relying on the SMPL-X. However, motion scanners imposed constraints on the environment, particularly in capturing large motions or events involving multiple humans. While AGORA [36] and ScoreHMR [46] sought to digress from the usage of the motion scanners by fitting human body models to the human motions in real-world images/videos, the quality of the fitted human models declined drastically on the images taken from a distance. One of our goals for our dataset was to incorporate multiple viewpoints, including distant views of the target scene. To avoid compromising the quality of human motions/poses, we chose to use motion capture devices, while implementing our approach to ensure that final results in the dataset are not limited by the environments in which the devices were used.
@@
Natural human motion acquisition. Whether we are creating a real or synthetic dataset, images of motions captured by directing humans to perform specific actions based on a description often appear awkward rather than natural. Because of that, most datasets aim to include humans engaged in daily activities (e.g., MS COCO [29], MPII Human Pose [3]), performing tasks such as sports (e.g., UCF-Sports [41], SoccerNet [11], SportsMOT [13]) or art (e.g., Human-Art [26]) to capture their motions and poses in the most natural states possible. However, it is self-contradictory to artificially create a virtual event to capture natural motions associated with the event. In this paper, we aim to detour this self-contradict by initially designing the virtual events (i.e., aforementioned games) using existing but non-natural virtual motions, which are then replaced by real-world motions captured using a motion capture device.
@@
Supplemental datasets for training. Enhancing model performance by supplementing the training with additional data has been a common strategy [38, 27]. Initially, this involved combining datasets constructed with the same purpose, like MS COCO [29] and PASCAL VOC [17], for tasks such as object detection. Some approaches utilized large-scale datasets (e.g., ImageNet [14] or Instagram [32]), which were not necessarily designed for the target task, to build foundational features, followed by a transfer learning such as pretrain-finetune [20] or PTL [43] to adjust the model on the target dataset. As models have grown in size and complexity (e.g., ViT [15]), the demand for large-scale, high-quality datasets has increased, but the high costs of annotation present a significant barrier. To address this, label-agnostic training methods like self-supervised learning [9, 8, 21, 22] and synthetic dataset generation with cost-free annotations have emerged as viable solutions. In response, we have specifically designed SynPlay to supplement various computer vision tasks that require a large-scale, highly-diversified human appearance set.
@@ 3 SynPlay Dataset
Diverse yet realistic human motions. We use rule-guided motion design in our SynPlay dataset, borrowing “rules” from six traditional Korean games, also featured in the Netflix series “Squid Game” [24] [^2] Our game scenarios are designed based on the traditional game rules, without taking any specific situations from the show.. This approach offers coarser motion guidance for the virtual players, facilitating the generation of a wide spectrum of natural motions, even including the ones that defy detailed description.
@@
Our rule-guided motion design is effectively baked into the overall sequence-generating design pipeline, as shown in Figure 2, that consists of the scenario design followed by the incorporation of real-world motions. The scenario design involves the setup of all the knobs that control the appearance, players (winners and losers), game dynamics (e.g., how/when each game ends), and the human motion evolutions for each specific scenario. All the items within each scenario that do not have to be hard-coded (e.g., game rules) are selected randomly when designing each scenario. The motion evolution of each virtual player in a specific game is governed by a graph structure where all possible elementary motions and their potential transitions are represented as nodes and directed edges, respectively. Each node is tied with a pool of motions that fall under the same elementary motion state (e.g., move, sit). As the game progresses, a virtual player evolves its motion following the directed edges and stays there according to the ‘motion timing’, also defined in the scenario design. At each state, the virtual human randomly chooses to exhibit one of the motions in the corresponding pool. Note that, while a uniquely designed scenario is used for a unique sequence, the same motion evolution graph is used for all the sequences captured under the same game rule.
@@
Before incorporating the real-world motions, we leverage two techniques to pre-diversify the elementary human motions readily available in human motion libraries such as Mixamo [2]: i) dynamically blending two existing motions of similar types to generate a new motion type (e.g., blending slow-walking and running to generate hasty-walking), ii) using elementary motions as animation layers to make a new motion (e.g., raising hands while walking). On top of the pre-diversified set of motions for each game, a real human player wearing a motion capture device  [^3] Each real player used a SmartSuit Pro II and a pair of Smartgloves from Rokoko (http://rokoko.com) is asked to either similarly mimic or newly create motions that align with the given game rule (thus, rule-guided). For example, for the game of tug-of-war, human players were provided with the game rule, and then asked to reenact any possible motion with the freedom of choosing the winning or losing side. For some games, more than one players were asked to actually play the game together to capture the motions that can naturally arise at the time of physical interactions. As the result of incorporating the real-world motions, the total number of unique motions in SynPlay increased from 104 to 257.
@@
UAV low-alt UAV med-alt UAV high-alt CCTV front CCTV side CCTV back UGV [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] Figure 3: Multiple viewpoints used in SynPlay. On the top-right corner of each image, we place the enlarged crop of one human instance who is visible from all seven viewpoints in each scenario. Multiple camera viewpoints allow substantial variations in appearance for the same human subject with identical pose.
Dataset specification. We created ten different scenarios for each game, resulting in 60 game scenarios in total. Frames were rendered from seven different camera viewpoints at 1 fps with the resolution of 1920$\times$1080, resulting in a total of 73,892 images with more than 6.5M human instances. The frame generation rate was selected as 1 fps to avoid including highly redundant human poses. Taking full advantage of the game engine’s ability to generate annotations while rendering the scenes, we provided various types of ground truth annotations useful for various computer vision tasks: 2D/3D bounding boxes, instance-level segmentation masks, depth maps, and human keypoint locations.

### v2

# SynPlay: Large-Scale Synthetic Human Data with Real-World Diversity for Aerial-View Perception

@@ Abstract
We introduce SynPlay, a large-scale synthetic human dataset purpose-built for advancing multi-perspective human localization, with a predominant focus on aerial-view perception. SynPlay departs from traditional synthetic datasets by addressing a critical but underexplored challenge: localizing humans in aerial scenes where subjects often occupy only tens of pixels in the image. In such scenarios, fine-grained details like facial features or textures become irrelevant, shifting the burden of recognition to human motion, behavior, and interactions. To meet this need, SynPlay implements a novel rule-guided motion generation framework that combines real-world motion capture with motion evolution graphs. This design enables human actions to evolve dynamically through high-level game rules rather than predefined scripts, resulting in effectively uncountable motion variations. Unlike existing synthetic datasets—which either focus on static visual traits or reuse a limited set of mocap-driven actions—SynPlay captures a wide spectrum of spontaneous behaviors, including complex interactions that naturally emerge from unscripted gameplay scenarios. SynPlay also introduces an extensive multi-camera setup that spans UAVs at random altitudes, CCTVs, and a freely roaming UGV, achieving true near-to-far perspective coverage in a single dataset. The majority of instances are captured from aerial viewpoints at varying scales, directly supporting the development of models for long-range human analysis—a setting where existing datasets fall short. Our data contains over 73k images and 6.5M human instances, with detailed annotations for detection, segmentation, and keypoint tasks. Extensive experiments demonstrate that training with SynPlay significantly improves human localization performance, especially in few-shot and data-scarce scenarios.
[Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image]
[Uncaptioned image]   [Uncaptioned image]   [Uncaptioned image]
Figure 1: SynPlay captures players performing six traditional games in a virtual playground, also featured in the Netflix series “Squid Game” [SquidGameNetflix2021]. It promotes motion diversity via rule-guided motion generation, where actions evolve dynamically through game mechanics. A multi-perspective setup enables near-to-far viewpoint coverage while capturing diverse behaviors and appearances across angles.
@@ 1 Introduction
Large-scale synthetic datasets have become a cornerstone in modern computer vision, providing scalable solutions for training large-capacity models [Unity, Unreal]. Advances in rendering engines such as Unity [Unity], Unreal Engine [Unreal], and human modeling tools like MakeHuman [MakeHuman] and Character Creator [Charactercreator] have enabled increasingly realistic representations of virtual humans in diverse environments. These tools have expanded the use of synthetic data across applications such as human pose estimation [PPatelCVPR2021, JLiaoCVPR2024], shape reconstruction [NMahmoodICCV2019], and semantic segmentation [SRichterECCV2016].
However, a critical but underexplored challenge remains: identifying humans from aerial perspectives where subjects occupy only tens of pixels. In these cases, detailed appearance cues such as facial features, textures, or fine-grained clothing differences become irrelevant. Instead, recognition must rely on factors like human motion, posture, and behavioral interactions. Despite the growing need for robust long-range human localization—relevant to tasks such as UAV-based surveillance, disaster response, and crowd monitoring—no existing dataset jointly addresses this challenge by providing both diverse human actions and extensive aerial viewpoint coverage.
Prior synthetic human datasets have primarily focused on enhancing diversity through individual traits, such as race, gender, and clothing [PPatelCVPR2021, MBlackCVPR2023], which are effective for close-range tasks like body/pose estimation or 3D reconstruction. However, their impact diminishes in distant-view scenarios where subjects are too small for these attributes to be resolved [NMahmoodICCV2019]. Moreover, datasets targeting pose diversity, such as SURREAL [GVarolCVPR2017] and AMASS [NMahmoodICCV2019], rely heavily on motion capture data obtained through physically constrained environments and predefined instructions (e.g., “5 seconds waving above the head with both arms” in HDM05 [NMahmoodICCV2019]). These approaches fail to capture the natural spontaneity of human behaviors, especially for complex multi-human interactions.
Existing aerial-view datasets, such as Okutama-Action [MBarekatainCVPRW2017], UAVid [YLyuPRS2020], VisDrone [PZhuTPAMI2022], and Archangel [YShenAccess2023], provide broader viewpoint coverage but suffer from limited motion diversity. Most prioritize capturing subjects from various altitudes and angles but recycle a small set of static or scripted motions. As a result, they do not reflect the breadth of real-world activities or emergent behaviors that occur in dynamic, unscripted scenarios. Synthetic datasets like SynDrone [GRizzoliICCVW2023] and CARGO [TZhangCVPR2021] similarly lack behavioral variability, typically limiting human actions to basic locomotion or pose transitions.
In this paper, we introduce SynPlay, a large-scale synthetic human dataset designed to close this gap. SynPlay departs from traditional synthetic data generation pipelines by introducing a rule-guided motion generation framework, where human motions evolve dynamically based on high-level game rules rather than predefined scripts. This framework combines real-world motion capture with motion evolution graphs, enabling a continuous expansion of behavior space. As a result, SynPlay provides effectively uncountable motion variations, capturing spontaneous human actions, including collisions, tugging, group interactions, and diverse game dynamics that naturally emerge in multi-human scenarios. For example, while datasets like Archangel-synth [YShenArXiv2024] or SynDrone [GRizzoliICCVW2023] contain only 2–3 distinct motions, SynPlay features a behaviorally diverse action space that cannot be easily enumerated (see Table 1 in the Supplementary material).
Furthermore, SynPlay incorporates a multi-perspective camera setup comprising three UAVs at randomized altitudes, three ground-based CCTVs at fixed viewpoints, and a freely roaming Unmanned Ground Vehicle (UGV). This design achieves true near-to-far perspective coverage within a single dataset, enabling robust localization of humans across extreme viewpoint variations. While some ground-level and close-up views are included, the majority of data consists of aerial observations where human subjects appear as small instances—often occupying only tens of pixels. This makes SynPlay uniquely suited for developing models that generalize across long-range aerial-view tasks, where prior datasets fall short.
To create SynPlay, we designed six traditional Korean games, also featured in the Netflix series “Squid Game” [SquidGameNetflix2021], as interactive scenarios. These games naturally elicit a broad range of dynamic human actions, including standing, sitting, grabbing, pulling, and wrestling, making them ideal for generating diverse behavior patterns in a controlled yet emergent manner. Each scenario is instantiated through a combination of motion blending, animation layering, and real-world motion capture, ensuring both realism and variability beyond the constraints of traditional mocap datasets.
Our main contributions are as follows:
  • We introduce SynPlay, the first large-scale synthetic human dataset specifically designed for identifying humans from aerial perspectives where subjects occupy only tens of pixels.
  • We propose a rule-guided motion generation framework, combining motion capture data with motion evolution graphs to create effectively uncountable motion variations, including spontaneous multi-human interactions.
  • We provide a multi-perspective data collection setup spanning UAVs, CCTVs, and UGVs, achieving true near-to-far viewpoint coverage within a single dataset.
  • We demonstrate through extensive experiments that SynPlay significantly improves aerial-view human detection and segmentation, especially in few-shot, cross-domain, and data-scarce scenarios.
1.1 Points to Note
SynPlay is synthetic data—but not the usual kind. Synthetic datasets offer essential benefits for vision tasks requiring large-scale data: effortless generation, cost-free labeling, and controllability. However, they often suffer from limited realism and motion diversity, especially in aerial-view human localization, where subjects occupy only a few dozen pixels in the image. Our core novelty addresses this by expanding the motion and behavior space, not just varying static traits. Prior aerial-view synthetic datasets typically focus on capturing multiple viewpoints while reusing a small set of motions, leading to repetitive behaviors seen from different angles. SynPlay resolves this by introducing rule-guided motion generation, where human actions evolve dynamically based on high-level game rules rather than predefined scripts. This framework combines real-world motion capture data with a motion evolution graph, enabling effectively uncountable motion variations. For example, while datasets like Archangel-synth or SynDrone contain only 2–3 distinct motions, SynPlay captures a wide spectrum of behaviors that naturally emerge from multi-human interactions. This includes spontaneous, unscripted actions that defy simple categorization. As detailed in the supplementary material (Tab. 1), this level of motion diversity is currently unmatched in aerial-view datasets.
SynPlay is primarily designed for multi-perspective human localization, with a predominant focus on aerial perception. SynPlay facilitates robust human localization across a broad range of perspectives and distances, with particular emphasis on aerial viewpoints. While it includes some ground-level and close-up views, the majority of the data consists of aerial observations where subjects often appear as only tens of pixels in size (see Fig. 2 in the supplementary material). This presents unique challenges distinct from tasks like pose recovery and mesh reconstruction, which rely on close-range, high-resolution imagery.
SynPlay is the only existing dataset to simultaneously achieve:
  • High instance-per-image density in aerial-view scenes (88.4 instances per image on average),
  • True near-to-far viewpoint coverage, including nadir-view observations,
  • Rule-guided human motion diversity, addressing the lack of behavioral variation in prior aerial-view datasets.
These characteristics make SynPlay a unique and valuable resource for advancing aerial-view human analysis, few-shot learning, and data-scarce cross-domain human localization tasks.
@@ 2 Related Works
Synthetic human data. The creation of various synthetic human datasets has been facilitated by the advancement of modern synthetic data rendering engines such as Blender, Unity, and Unreal, alongside human modeling tools like MakeHuman [MakeHuman] and Character Creator [Charactercreator]. These rendering engines enable a realistic representation of humans in 3D virtual environments, while the modeling tools give creators precise control over the design of virtual characters. The creators of these datasets leveraged these tools to meticulously control key design factors, ensuring suitability for specific tasks, e.g., SOMAset [IBarbosaCVIU2018], PersonX [XSunCVPR2019], UnrealPerson [TZhangCVPR2021], CARGO [QZhangCVPR2024] for Re-Identification, SURREAL [GVarolCVPR2017] for pose estimation, GTA5 [SRichterECCV2016] for semantic segmentation, and Archangel-Synthetic [YShenAccess2023] for detection.
Recently, various efforts have aimed to enhance the realism of virtual human models, bringing them closer to real-world quality. SMPL-X [GPavlakosCVPR2019], AMASS [NMahmoodICCV2019], and HardMo [JLiaoCVPR2024] leveraged motion capture devices to record natural human motions. BEDLAM [MBlackCVPR2023] expanded diversity in factors like skin tones and clothing, though still based on SMPL-X. However, motion capture systems impose environmental constraints, especially when capturing large-scale motions or multi-human interactions. AGORA [PPatelCVPR2021] and ScoreHMR [AStathopoulosCVPR2024] attempted to bypass motion scanners by fitting body models to real-world images and videos, but model quality degrades significantly for distant-view images. A key goal of our dataset was to include multiple viewpoints, including distant perspectives, without sacrificing motion fidelity. To this end, we used motion capture devices but designed our pipeline to ensure the dataset is not limited by the recording environments.
@@
Natural human motion acquisition. Whether we are creating a real or synthetic dataset, images of motions captured by directing humans to perform specific actions based on a description often appear awkward rather than natural. Because of that, most datasets aim to include humans engaged in daily activities (e.g., MS COCO [TLinECCV2014], MPII Human Pose [MAndrilukaCVPR2014]), performing tasks such as sports (e.g., UCF-Sports [MRodriguezCVPR2008], SoccerNet [ACioppaCVPRW2022], SportsMOT [YCuiICCV2023], PKU-DyMVHumans [XZhengCVPR2024], SportsHHI [TWuCVPR2024]) or art (e.g., Human-Art [XJuCVPR2023]) to capture their motions and poses in the most natural states possible. However, it is self-contradictory to artificially create a virtual event to capture natural motions associated with the event. In this paper, we aim to detour this self-contradict by initially designing the virtual events (i.e., aforementioned games) using existing but non-natural virtual motions, which are then replaced by real-world motions captured using a motion capture device.
@@
Supplemental datasets for training. Enhancing model performance by supplementing the training with additional data has been a common strategy [SRenTPAMI2016, HLeeTIP2019]. Initially, this involved combining datasets constructed with the same purpose, like MS COCO [TLinECCV2014] and PASCAL VOC [MEveringhamIJCV2015], for tasks such as object detection. Some approaches utilized large-scale datasets (e.g., ImageNet [JDengCVPR2009] or Instagram [DMahajanECCV2018]), which were not necessarily designed for the target task, to build foundational features, followed by a transfer learning such as pretrain-finetune [RGirshickTPAMI2016] or PTL [YShenCVPR2023] to adjust the model on the target dataset. As models have grown in size and complexity (e.g., ViT [ADosovitskiyICLR2021]), the demand for large-scale, high-quality datasets has increased, but the high costs of annotation present a significant barrier. To address this, label-agnostic training methods like self-supervised learning [TChenICML2020, MCaronNeurIPS2020, JGrillNeurIPS2020, KHeCVPR2022] and synthetic dataset generation with cost-free annotations have emerged as viable solutions. In response, we have specifically designed SynPlay to supplement various computer vision tasks that require a large-scale, highly-diversified human appearance set.
@@ 3 SynPlay Dataset
Diverse yet realistic human motions. We use rule-guided motion design in our SynPlay dataset, borrowing “rules” from six traditional Korean games, also featured in the Netflix series “Squid Game” [SquidGameNetflix2021] [^1] Our game scenarios are designed based on the traditional game rules, without taking any specific situations from the show.. This approach offers coarser motion guidance for the virtual players, facilitating the generation of a wide spectrum of natural motions, even including the ones that defy detailed description.
@@
Our rule-guided motion design is effectively baked into the overall sequence generation pipeline, as shown in Figure 2, that consists of the scenario design followed by the incorporation of real-world motions. The scenario design involves the setup of all the knobs that control the appearance, players (winners and losers), game dynamics (e.g., how/when each game ends), and the human motion evolutions for each specific scenario. In each scenario, all items that don’t need to be hard-coded (such as game rules) are selected randomly during scenario design. The motion evolution of each virtual player in a specific game is governed by a graph structure where all possible elementary motions and their potential transitions are represented as nodes and directed edges, respectively. Each node is tied with a pool of motions that fall under the same elementary motion state (e.g., move, sit). As the game progresses, a virtual player evolves its motion following the directed edges and stays at a node according to the ‘motion timing’, also defined in the scenario design. In each state, the virtual human randomly chooses to exhibit one of the motions in the corresponding pool. It is noteworthy to mention that, while our motion evolution graphs are rule-guided, they are compositional and parameterized. This yields a combinatorial space, effectively generating unbounded variants rather than a fixed library.
@@
Before incorporating the real-world motions, we leverage two techniques to pre-diversify the elementary human motions readily available in human motion libraries such as Mixamo [Mixamo]: i) dynamically blending two existing motions of similar types to generate a new motion type (e.g., blending slow-walking and running to generate hasty-walking), ii) using elementary motions as animation layers to make a new motion (e.g., raising hands while walking). On top of the pre-diversified set of motions for each game, a real human player wearing a motion capture device  [^2] Each real player used a SmartSuit Pro II and a pair of SmartGloves from Rokoko (http://rokoko.com) is asked to either similarly mimic or newly create motions that align with the given game rule (thus, rule-guided). For example, for the game of tug-of-war, human players were provided with the game rule, and then asked to reenact any possible motion with the freedom of choosing the winning or losing side. For some games, more than one players were asked to actually play the game together to capture the motions that can naturally arise at the time of physical interactions. As the result of incorporating the real-world motions, the total number of unique motions in SynPlay increased from 104 to 257.
UAV low-alt UAV med-alt UAV high-alt CCTV front CCTV side CCTV back UGV
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: Multiple viewpoints used in SynPlay. Multiple camera viewpoints allow substantial variations in appearance for the same human subject with identical pose.
Multiple viewpoints. The camera viewpoints within SynPlay are diversified by implementing three widely used types of image-capturing platforms in the real world: UAV, UGV, and CCTV. They cover a variety of image-capturing properties like static/dynamic and ground/aerial. Viewpoint diversity is acquired by controlling the locations and the focal points of the cameras. Three UAVs, three CCTVs, and one UGV have been deployed (Figure 3), resulting in seven unique viewpoints for every game sequence. The UAVs are deployed to fly at various random locations while maintaining altitudes of low ($\sim$30mm), medium ($\sim$50mm), and high ($\sim$100mm). CCTVs are located at a height of 15mm at the front, back, and either side of the game playground. UGV images are captured assuming that a vehicle is randomly roaming on the ground. The focal points are set at several locations close to the area where the game usually takes place. For the UAVs, the focal point is changed to a random location every 10 $\mathit{sec}$, where each change takes 5 $\mathit{sec}$ to be fixed at a location for another 5 $\mathit{sec}$. Focal points of the CCTVs and the UGV do not change once determined.